How do you find the area for this expression?: Erik has 2000 yards of fencing to enclose a rectangular area. Find the dimensions of the rectangle that maximize the enclosed area. What is the maximum area?
Answer by stanbon(75887)

How do you find the area for this expression?: Erik has 2000 yards of fencing to enclose a rectangular area. Find the dimensions of the rectangle that maximize the enclosed area. What is the maximum area?
----------------
Perimeter = 2(length + width)
2000 = 2(length + width)
length + width = 1000
L = 1000-W
----------------------
Area = Length*Width
A = (1000-W)*W
----
Area = 1000W-W^2
------
Maximum Area occurs when W = -b/(2a) = -1000(2*-1) = 500
-----
If Width = 500, Length = 1000 - 500 = 500
-----------------
Ans: Area = 500^2 = 250000 sq. units
--------------
Cheers,
Stan H.